OpenStudy (amistre64):

If x^2+x = 3 then x^4+x = ??

OpenStudy (anonymous):

whats the idea behind ur question? :)

OpenStudy (amistre64):

i got an old GRE math practice test ;)

OpenStudy (amistre64):

im thinking let x = 3-x^2 and let x^2 = 3-x and sub into the other one

OpenStudy (amistre64):

options are: A. x^2+3 B. 4(3-x) C. 3(x-1) D. 6(2-x) E. 3x i havent checked the answer key tho ...

OpenStudy (anonymous):

\[3-x^2+(x-3)^2=12-6x=6(2-x)\]unfortunately GRE test is too expensive in my country :(

OpenStudy (anonymous):

so D :)

OpenStudy (amistre64):

i like D answer key likes D tehn D it is :)